### v1

# A kinetic theory approach to consensus formation in financial markets

Abstract.
We investigate the relationship between analysts’ one-year consensus forecasts for the S&P 500 index and its current level. Contrary to the conventional view that sentiment drives price forecasts, our analysis predicts that the average consensus has no effect on forecasts, while the current S&P 500 index level alone is sufficient to anticipate analysts’ price expectations.
Employing a kinetic theory framework, we model the dynamics of analysts’ opinions, by taking into account both the mutual influences shaping price consensus and the dynamics of the actual S&P 500 index level. Testing the model on real data on a long-time series of data shows that just three free parameters are enough to accurately describe the one-year average price forecasts of analysts.
1. Introduction
Opinion dynamics play an important role in understanding the emergence of collective behaviors in markets, particularly under external influences such as price movements. Conventional perspectives often link price evolution to shifts in market sentiment. In contrast, our work investigates the reverse relationship: how market sentiment reacts to and closely follows price dynamics. This approach is essential for unraveling the feedback loops that drive decision-making within financial systems.
The application of methods from statistical mechanics to model social and economic systems has a long and successful history. This interdisciplinary approach was pioneered by Weidlich [15] and further popularized by Galam, Gefen, and Shapir [6]. By treating individuals as interacting entities and modeling their collective behavior, researchers have gained insights into complex social phenomena. In particular, kinetic theory has emerged as a powerful tool for describing the evolution of distributions in such systems, especially in the case of continuous opinion variables.
Indeed, kinetic theory provides a natural framework for modeling the consensus formation process in an interacting community. Originally developed in statistical physics to describe the evolution of particle distributions in a gas, kinetic models have been successfully applied to various domains, including opinion dynamics and economic interactions. These models typically describe the time evolution of a probability distribution, capturing the stochastic nature of agent interactions and the emergence of macroscopic patterns from microscopic rules (see, for example, [1, 2, 3, 8, 9, 10, 14]).
In the context of financial forecasting, analysts continuously update their expectations based, for example, on new information, market trends, and (possibly) peer influence. This process can be viewed as an interaction-driven system, where opinions evolve through repeated exchanges and adjustments. Unlike deterministic approaches that yield a single forecasted value, kinetic models allow for a richer description by capturing the full distribution of expectations, thus accounting for heterogeneity in analysts’ predictions. Furthermore, such models can describe key phenomena such as polarization, clustering of forecasts, and the convergence toward consensus.
A kinetic approach is particularly well suited for studying the expected level of the S&P 500 index at a one-year horizon, as it allows us to incorporate both endogenous factors, such as analysts’ interactions, and exogenous influences, including macroeconomic shocks and financial news. By extending methods from kinetic theory to this setting, we aim to provide a realistic and flexible model for predicting market sentiment, ultimately improving our understanding of how collective financial expectations evolve over time.
Specifically, we focus on the one-year consensus forecasts of analysts aggregated across all stocks belonging in the S&\&&P 500 index. We first demonstrate that the S&\&&P 500 and its consensus target price, more precisely, the logarithms of these two series, are cointegrated. This property is easily confirmed by using Engle-Granger procedure (see [5]). The short-run dynamics is also investigated thanks to Johansen procedure (see [11] and [12]). While it is not unexpected that these two series share a common trend, we find, surprisingly, that Granger causality test leads to the conclusion that the S&\&&P 500 level causes the consensus target price, but not vice versa. This result appears to be contradictory with several studies on the impact of analyst revisions on prices (see for instance [4], [13]). It is important to note, however, that these authors focus on analyst earnings revisions, while we examine their ability to forecast prices at a one-year horizon. Furthermore, these studies focus on individual stocks, while we are interested in the aggregate forecast for the entire S&\&&P 500.
Consequently, a central hypothesis of our study, while designing the kinetic model, is that the consensus does not predict or influence market movements, but rather follows them. Our model assumes that two main factors drive the analysts’ opinions: the influence of other analysts (described as a mean-field interaction) and the actual price of the index. This framework allows us to explore the interplay between these two effects and their impact on opinion formation.
Furthermore, our analysis shows that the influence of the mean-field becomes negligible over time, leaving the current price as the sole determinant of the average one-year forecasts. Hence, the average of the analysts’ one-year forecasts on the S&P 500 index can be explained using only three fixed parameters: $\Delta$, representing a percentage of the current price; $q$, a coefficient for the convex combination of the current index value and the analyst’s current opinion; and $\beta$, the relative weight of following the index value versus the mean-field.
By demonstrating the reactive nature of the consensus forecast and quantifying their dependence on price movements, this work contributes to a deeper understanding of opinion dynamics in financial markets.
2. exploratory data analysis
2.1. Datas
Our study is based on the one-year target price published by Bloomberg for the S&P 500 across all analysts. More specifically, bank or broker analysts publish their one-year target prices for equities, particularly those in the S&P 500. Bloomberg collects and distributes these datas to market participants via its terminal. Then, it calculates and publishes a one-year consensus forecast for each stock based on the median of the whole one-year target prices. Bloomberg finally computes the sum of one-year consensus for S&P 500 stocks weighted by the weight of each stock in the index to obtain a one-year consensus forecast for the level of the S&P 500 based on analyst recommendations. The daily data analyzed in this article cover the period from May 26th, 2009, to November 20th, 2024.
Figure 1. Time series of the S&P 500 index $X(t)$ and the one-year analysts’ forecasts for the S&P 500 index, from May 26th, 2009, to November 20th, 2024. Source: Bloomberg
In the remainder of this paper, $X(t)$ will denote the level of the S&P 500 at time $t$ while $s(t)$ will be the one year consensus target price computed by Bloomberg for the S&P 500.
2.2. Long-run relationship between log price and log target price for S&P 500
It is quite obvious that $\ln(X(t))$ and $\ln(s(t))$ are nonstationary and, indeed, stationarity is rejected by the Augmented Dickey-Fuller test at any confidence level.
For a discrete time series $Y(t)$, define the first difference $DY(t)=Y(t)-Y(t-1)$. This time, $D\ln(X(t))$ and $D\ln(s(t))$ are accepted to be stationary according to the ADF test (at $1\%$ confidence level). The time series $\ln(X(t))$ and $\ln(s(t))$ are then said to be $I(1)$ (integrated of order 1) and their first differences, which are stationary time series, are said to be $I(0)$ (integrated of order 0).
Now, we investigate the cointegration between $\ln(X(t))$ and $\ln(s(t))$, which is a linear relation linking theses series. In that particular case, the cointegration relationship would be defined as follows:
Definition 2.1.
Suppose discrete time series $Y(t)$ and $Z(t)$ are $I(1)$. Then $Y(t)$ and $Z(t)$ are said to be cointegrated (of order 1) if there exist $\alpha$, $\beta$ and a time series $u(t)$ which is $I(0)$ such that:
(1) $Y(t)=\alpha Z(t)+\beta+u(t).$
Remark 2.2.
Concepts of integration and cointegration can easily be generalized to higher order but we don’t need this general framework here.
Usually, one cannot use the Ordinary Least Square method to estimate parameters in the cointegration relationship (1) for it leads to false conclusions. This issue has first been exhibited by Yule (see [16]). Decades later, Granger and Newbold investigated this misprocedure more deeply [7] and baptised it “spurious regression”. Thereafter, Engle &\&& Granger ([5]) proposed the following procedure to investigate cointegration and estimate the parameters whenever $Y(t)$ and $Z(t)$ are proved to be $I(1)$:
    First, estimate the (long-run) relationship (1) between $Y$ and $Z$ by computing OLS for the linear regression of $Y$ on $Z$ (even though it is not allowed).
    Then compute $e(t)=Y(t)-\widehat{\alpha}Z(t)-\widehat{\beta}$.
    At last, test stationarity for $e(t)$ by using Augmented-Dickey Fuller test. If stationarity is accepted, then $Y(t)$ and $Z(t)$ are cointegrated and $\widehat{\alpha}$ and $\widehat{\beta}$ are convergent estimators for $\alpha$ and $\beta$.
The estimation of the linear regression $\ln(s(t))=\alpha\ln(X(t))+\beta+u(t)$ leads to $\widehat{\alpha}=0.98679$ and $\widehat{\beta}=0.21778$.
Augmented Dickey-Fuller test allows us to accept stationarity for $e(t)=\ln(s(t))-0.98679\ln(X(t))-0.21778$ (at $1\%$ confidence level).
This leads to the conclusion that $\ln(X(t))$ and $\ln(s(t))$ are cointegrated. This appears clearly by looking at figure 2 which compares $\ln(s(t))$ with its estimated long-term equilibrium level $\alpha\ln(X(t))+\beta$.
Figure 2. Logarithm of one-year analysts’ forecasts for the S&P 500 index vs its long-term equilibrium level, from May 26th, 2009, to November 20th, 2024.
2.3. Short-run relationship
One major result regarding cointegration theory is the Granger-Engle Representation Theorem (see [5]):
Theorem 2.3.
Suppose $Y(t)$ and $Z(t)$ are cointegrated at the order 1. Then there exists an error-correction representation i.e. an (error-correction) model of the form:
(2) $DY(t)=\sum_{i=1}^{p}a_{i}DY(t-i)+\sum_{j=1}^{p}b_{j}DZ(t-j)+\gamma\left(Y(t-1)-\alpha Z(t-1)-\beta\right)+\varepsilon(t),$
where $\varepsilon(t)$ is $I(0)$.
Remark 2.4.
The name of Model (2) stems from the fact that the error in the cointegration relationship at time $t-1$ is corrected at time $t$ by the restoring force $\gamma$, which must be negative. The more the “error” at time $t-1$ in the cointegration relationship is large, the stronger the restoring force is.
One can notice that there also exists an error correction model using $DZ(t)$ as the dependent variable since Equation (1) can be written:
$Z(t)=\frac{1}{\alpha}Y(t)-\frac{\beta}{\alpha}-\frac{1}{\alpha}u(t).$
Hence, we have to investigate two short-term equations, but we keep the same long-run relationship in both models, $s(t)$ being the response variable. Consequently, $\gamma^{\prime}$ estimated below will remain negative while $\gamma$ will become positive. Based on three out of four information criteria (Akaike, Schwarz, Hannan-Quin and Final Prediction error), the optimal number of lags $p$ for short-run dynamics is estimated to be 9 (see Table 1). This is therefore the value that will be retained to estimate the short-term equations.
AIC SIC HQ FPE
9 5 9 9
Table 1. Optimal number of lags for each criterion
2.3.1. Short-run relationship with $D\ln(X(t))$ as the dependent variable
The first model we look at is:
$D\ln\left(X(t)\right)=\sum_{i=1}^{9}a_{i}D\ln(X(t-i))+\sum_{j=1}^{9}b_{j}D\ln(s(t-j))+\gamma\left(\ln(s(t-1))-\alpha\ln(X(t-1))-\beta\right)+\varepsilon(t).$
The model is estimated in R according to Johansen procedure by using cajorls and vecm functions. None of the coefficients of the past level of $D\ln(s(t))$ (i.e. from $b_{1}$ to $b_{9}$) are significant (see Table 2 below). Hence $\ln(s(t))$ does not (Granger-)cause $\ln(X(t))$. The $R^{2}$ is only $2\%$, which is very small. We can deduce that the short-term dynamics of the market is not impacted at all by the consensus. One can remark, though, that there is a bit of autocorrelation for $\ln(X(t))$ (e.g. $a_{1}$ is significant).
Parameter Estimate Std $t$-stat
$\gamma$ 0.002183 0.005223 0.418
$a_{1}$0.087234 0.015780 -5.528***
$a_{2}$ 0.053540 0.015902 3.367***
$a_{3}$0.028266 0.015948 -1.772 $\cdot$
$a_{4}$0.025171 0.015985 -1.575
$a_{5}$0.010029 0.016015 -0.626
$a_{6}$0.063941 0.016094 -3.973***
$a_{7}$ 0.054246 0.016146 3.360***
$a_{8}$0.044517 0.016156 -2.755**
$a_{9}$ 0.060541 0.016119 3.756***
$b_{1}$0.106482 0.208119 -0.512
$b_{2}$ 0.190598 0.220288 0.865
$b_{3}$ 0.179068 0.220953 0.810
$b_{4}$0.024094 0.221821 -0.109
$b_{5}$ 0.226850 0.221784 1.023
$b_{6}$0.218241 0.221501 -0.985
$b_{7}$ 0.131186 0.220381 0.595
$b_{8}$0.266791 0.219320 -1.216
$b_{9}$0.138858 0.205359 -0.676
Table 2. Estimates of the short-term equation for the response variable $D\ln(X(t))$
***, **, *, $\cdot$ indicate stastistical significance at $0.1\%$, $1\%$, $5\%$, $10\%$ levels respectively
2.3.2. Short-run relationship with $D\ln(s(t))$ as the dependent variable
The second model is written:
$D\ln(s(t))=\sum_{i=1}^{9}a^{\prime}_{i}D\ln(s(t-i))+\sum_{j=1}^{9}b^{\prime}_{j}D\ln(X(t-j))+\gamma^{\prime}\left(\ln(s(t-1))-\alpha\ln(X(t-1))-\beta\right)+\varepsilon^{\prime}(t).$
The short-term equation with $D\ln(s(t))$ as the dependent variable is reported in Table 3. The $R^{2}$ of the model is around $60\%$, wich is quite good, and most of the coefficients are significant.
Parameter Estimate Std $t$-stat
$\gamma^{\prime}$0.0047212 0.0003967 -11.901***
$a^{\prime}_{1}$ 0.3476735 0.0158070 21.995 ***
$a^{\prime}_{2}$ 0.0829855 0.0167312 4.960 ***
$a^{\prime}_{3}$ 0.0978602 0.0167817 5.831 ***
$a^{\prime}_{4}$ 0.0643269 0.0168477 3.818 ***
$a^{\prime}_{5}$ 0.0582016 0.0168449 3.455 ***
$a^{\prime}_{6}$ 0.0195366 0.0168233 1.161
$a^{\prime}_{7}$ 0.0014429 0.0167383 0.086
$a^{\prime}_{8}$0.0271617 0.0166577 -1.631
$a^{\prime}_{9}$ 0.0157422 0.0155973 1.009
$b^{\prime}_{1}$ 0.0074338 0.0011985 6.202 ***
$b^{\prime}_{2}$ 0.0063950 0.0012078 5.295 ***
$b^{\prime}_{3}$ 0.0066243 0.0012113 5.469 ***
$b^{\prime}_{4}$ 0.0050619 0.0012141 4.169 ***
$b^{\prime}_{5}$ 0.0080997 0.0012164 6.659 ***
$b^{\prime}_{6}$ 0.0045694 0.0012224 3.738 ***
$b^{\prime}_{7}$ 0.0043499 0.0012263 3.547 ***
$b^{\prime}_{8}$ 0.0053319 0.0012271 4.345 ***
$b^{\prime}_{9}$ 0.0061260 0.0012243 5.004 ***
Table 3. Estimates of the short-term equation for the response variable $D\ln(s(t))$
***, **, *, $\cdot$ indicate stastistical significance at $0.1\%$, $1\%$, $5\%$, $10\%$ levels respectively
This equation is the most relevant one, as it demonstrates that the consensus target price for the S&\&&P 500 adjusts based on the market level, rather than the other way around. As a result, the consensus acts more as a “market follower” than a “market mover” in the short term.
Remark 2.5.
One can remark that $\alpha$ in the long-run equation is near 1. The Augmented Dickey-Fuller test for $\ln(s(t))-\ln(X(t))$ leads to the conclusion that one can accept that $\alpha$ is equal to 1, which, again, is no surprise when looking at Figure 1.
Remark 2.6.
The half-life of the second short-run relationship is the number of periods that are required to correct half the deviation from equilibrium in the long-run relationship. It is computed by $t$ such that
$(1+\gamma^{\prime})^{t}=\frac{1}{2}$
which yields to $t=147$ trading days, i.e. between 6 and 7 months.
All these results permits us to design our kinetic model which, in return, will improve our understanding of the long-term dynamics linking the consensus to the market.
3. The kinetic model
The kinetic-type model described in this Section predicts the behavior of the S&P 500 index by supposing that the influence on the index forecasts is governed by a small number of inputs, namely the opinions of the other agents (consensus), the market suggestions, and the degree of attractiveness of the share price.
The model’s independent variables are time, $t\in\mathbb{R}^{+}$, and the one-year price forecast for a given agent, denoted by $x\in\mathbb{R}_{+}$.
The agents’ community is described by the distribution function $f=f(t,x)$, defined on +×$\mathbb{R}^{+}\times\bar{\Omega}$. This function represents the density of agents holding price forecast $x$ at time $t$. Its evolution is governed by a kinetic equation, encompassing two levels of interaction: the microscopic level, where opinions on the price forecasts are influenced by the mean field and external market forces, and the macroscopic level, which captures the evolution of $f(t,x)$ induced by these microscopic interactions.
Interaction with the price dynamics
The influence of price dynamics on the one-year price forecast evolution is governed by four primary features:
    Strength of the consensus on the forecast. It is described by a time-dependent parameter $\alpha\in\mathbb{R}^{+}$ that quantifies the intensity of the influence exerted by the forecast of the other agents on the evolution of one-year price forecasts.
    The share price. It is described by a function $X(t):\mathbb{R}^{+}\to\bar{\Omega}$. The perfect knowledge of the share price reflects changes in the market environment and serves as the primary external factor affecting individual opinions on price forecasts.
    Attractiveness. The attractiveness is described by a parameter $q\in(0,1)$ that quantifies the degree to which a one-year price forecast $x$ is drawn toward the share price $X(t)$.
    Tendency to consensus. This feature implies that opinions depend not only on the attractiveness of the share price at time $t$, but also on the mean field of price opinions.
The central mechanism of the model lies in the interaction rule, which reflects the natural tendency of opinions to converge while being influenced by price dynamics. Following [3], we describe the evolution of opinions as:
(3) $\bar{x}=(1-q)x+qX(t)(1+\Delta),$
where $\Delta\in\mathbb{R}$ is a spread function that encapsulates the deviation of forecasts from the actual market price.
We assume that the time dynamics of $X(t)$ is known. For technical reasons, we suppose moreover that the share price is a continuous and uniformly bounded function of time.
Inverting the interaction rule in (3), we obtain:
$x=\frac{1}{1-q}\bar{x}-\frac{q}{1-q}(X(t)+\Delta).$
Time Evolution of $f$
We first introduce the sentiment of the population on the target price $s\,:\,\mathbb{R}^{*}_{+}\to\mathbb{R}_{+}$, defined as:
(4) $s(t)=\frac{\displaystyle\int_{\mathbb{R}_{+}}xf(t,x)\,\mathrm{d}x}{\displaystyle\int_{\mathbb{R}_{+}}f(t,x)\,\mathrm{d}x}.$
The time evolution of the opinion distribution $f$ is governed by a non-local partial differential equation. In its classical form, the model is the following:
(5) $\left\{\begin{array}[]{l}\displaystyle\partial_{t}f+\alpha\partial_{x}\big{(}(s(t)-x)f\big{)}=\displaystyle\beta\left[\frac{1}{1-q}f\left(\frac{x}{1-q}-\frac{q}{1-q}X(t)(1+\Delta)\right)\mathds{1}_{\left(\frac{x}{1-q}-\frac{q}{1-q}X(t)(1+\Delta)\geq 0\right)}-f(x)\right]\\[10.0pt] f(0,x)=f^{\text{in}}(x)\qquad x\in{\mathbb{R}^{*}_{+}}\\[10.0pt] f(t,0)=0\qquad t\in{\mathbb{R}^{*}_{+}},\\[10.0pt] \end{array}\right.$
where $\alpha\in\mathbb{R}_{+}^{*}$ is a parameter which describes how strong is the comparison with the sentiment of the population and $\beta\in\mathbb{R}_{+}^{*}$ is the characteristic relaxation time to the spot price.
The weak formulation
It is convenient to express and to study the previous model in weak form. Let $\phi\in C_{c}^{1}(\mathbb{R}\times\mathbb{R})$, which is the space of funtions of class $C^{1}$ with compact support in $\mathbb{R}\times\mathbb{R}$. From (5), we deduce formally that
(6) $\begin{array}[]{l}\displaystyle\frac{\mathrm{d}}{{\rm d}t}\int_{\mathbb{R}_{+}}f(t,x)\phi(t,x){\rm d}x+\alpha\int_{\mathbb{R}_{+}}\big{(}(s(t)-x)f\big{)}\partial_{x}\phi(t,x){\rm d}x=\displaystyle\beta\int_{\mathbb{R}_{+}}f(t,x)\left[\phi(t,\bar{x})-\phi(t,x)\right]{\rm d}x\end{array}$
for all $\phi\in C_{c}^{1}(\mathbb{R}\times\mathbb{R})$.
Denote
$m_{0}(t)=\int_{\mathbb{R}_{+}}f(t,x){\rm d}x\ \text{ and }\ m_{1}(t)=\int_{\mathbb{R}_{+}}xf(t,x){\rm d}x.$
When $\phi=1$, we deduce from the weak form that
$\frac{\mathrm{d}}{{\rm d}t}m_{0}=0,$
i.e.
$m_{0}=\int_{\mathbb{R}_{+}}f^{\text{in}}(x){\rm d}x$
for all $t\in\mathbb{R}_{+}$.
On the other hand, when $\phi=x$, the weak form (6) reduces to the following ordinary differential equation governing the time evolution of $m_{1}$:
$\frac{\mathrm{d}}{{\rm d}t}m_{1}(t)=-q\beta m_{1}(t)+q\beta X(t)(1+\Delta)m_{0}.$
If we divide by the constant $m_{0}$, we deduce an ordinary differential equation which describes the time evolution of the sentiment of the population:
$\frac{\mathrm{d}}{{\rm d}t}s(t)=q\beta\big{[}X(t)(1+\Delta)-s(t)\big{]}.$
This equation can be explicitly solved. Its solution is:
(7) $s(t)=\left(\int_{\mathbb{R}_{+}}xf^{\text{in}}(x)\mathrm{d}x\right)\left(\int_{\mathbb{R}_{+}}f^{\text{in}}(x)\mathrm{d}x\right)^{-1}e^{-q\beta t}+q\beta\int_{0}^{t}X(s)(1+\Delta)e^{q\beta(s-t)}\mathrm{d}s.$
Hence, from now on, Equation (7) will replace Equation (4) within the initial-boundary value problem (5).
4. Existence and uniqueness of the solution
By developing the transport term, our non-local partial differential equation can be written in the following form:
$\partial_{t}f+\alpha\left(s(t)-x\right)\partial_{x}f=\frac{\beta}{1-q}f\left(\frac{x}{1-q}-\frac{q}{1-q}X(t)(1+\Delta)\right)\mathds{1}_{\left(\frac{x}{1-q}-\frac{q}{1-q}X(t)(1+\Delta)\geq 0\right)}+(\alpha-\beta)f(t,x).$
Thanks to the method of characteristics, the previous equation can be written in integral form:
(8) $\begin{array}[]{l}\displaystyle f(t,x)=f^{\text{in}}\left(xe^{\alpha t}-\alpha\int_{0}^{t}s(\theta)e^{\alpha\theta}\mathrm{d}\theta\right)e^{(\alpha-\beta)t}\mathds{1}_{\left[\left(xe^{\alpha t}-\alpha\int_{0}^{t}s(\theta)e^{\alpha\theta}\mathrm{d}\theta\right)\geq 0\right]}+\\[10.0pt] \displaystyle\frac{\beta}{1-q}\int_{0}^{t}e^{(\beta-\alpha)(\theta-t)}f\left(\frac{x}{1-q}-\frac{q}{1-q}X(\theta)(1+\Delta)\right)\mathds{1}_{\left(\frac{x}{1-q}-\frac{q}{1-q}X(\theta)(1+\Delta)\geq 0\right)}\mathrm{d}\theta,\end{array}$
with $s$ given by (7).
Let $T\in\mathbb{R}_{+}^{*}$ and $C_{b}([0,T]\times\mathbb{R}_{+})$ the space of uniformly bounded continuous functions on $[0,T]\times\mathbb{R}_{+}$, such that
$\|f\|_{C_{b}([0,T]\times\mathbb{R}_{+})}=\max_{(t,x)\in[0,T]\times\mathbb{R}_{+}}|f(t,x)|<+\infty.$
The following result holds:
Theorem 4.1.
The integral problem (8), with initial condition $f^{\text{in}}\in C_{b}(\mathbb{R}_{+})$, has one and only one solution $f\in C_{b}([0,T]\times\mathbb{R}_{+})$.
Proof.
Denote with
$F(f^{\text{in}})=f^{\text{in}}\left(xe^{\alpha t}-\alpha\int_{0}^{t}s(\theta)e^{\alpha\theta}\mathrm{d}\theta\right)e^{(\alpha-\beta)t}\mathds{1}_{\left[\left(xe^{\alpha t}-\alpha\int_{0}^{t}s(\theta)e^{\alpha\theta}\mathrm{d}\theta\right)\geq 0\right]}$
and with $\mathscr{T}\,:\,C_{b}([0,T]\times\mathbb{R}_{+})\to C_{b}([0,T]\times\mathbb{R}_{+})$ the operator such that
$\mathscr{T}g=\frac{\beta}{1-q}\int_{0}^{t}e^{(\beta-\alpha)(\theta-t)}g\left(\frac{x}{1-q}-\frac{q}{1-q}X(\theta)(1+\Delta)\right)\mathds{1}_{\left(\frac{x}{1-q}-\frac{q}{1-q}X(\theta)(1+\Delta)\geq 0\right)}\mathrm{d}\theta,$
where $g\in C_{b}([0,T]\times\mathbb{R}_{+})$.
Hence, Equation (8) can be written as a fixed-point problem:
(9) $f=F(f^{\text{in}})+\mathscr{T}f$
in $C_{b}([0,T]\times\mathbb{R}_{+})$. In what follows, we show that (9) has one and only one solution in $C_{b}([0,T]\times\mathbb{R}_{+})$.
We first notice that
$\|F(f^{\text{in}})\|_{C_{b}([0,T]\times\mathbb{R}_{+})}\leq\|f^{\text{in}}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}e^{|\alpha-\beta|T}.$
On the other hand, the operator $\mathscr{T}$ is clearly linear and is bounded in $C_{b}([0,T]\times\mathbb{R}_{+})$ with bounded norm $\|\mathscr{T}\|_{\mathcal{L}(C_{b}([0,T]\times\mathbb{R}_{+}))}$. Indeed,
$\|\mathscr{T}g\|_{C_{b}([0,T]\times\mathbb{R}_{+})}\leq\|g\|_{C_{b}([0,T]\times\mathbb{R}_{+})}\left(\frac{\beta}{1-q}\right)\left(\frac{1-e^{-(\beta-\alpha)T}}{\beta-\alpha}\right)$
for all $g\in C_{b}([0,T]\times\mathbb{R}_{+})$.
Hence,
$\|\mathscr{T}\|_{\mathcal{L}(C_{b}([0,T]\times\mathbb{R}_{+}))}=\sup_{\|g\|_{C_{b}([0,T]\times\mathbb{R}_{+})}=1}\|\mathscr{T}g\|_{C_{b}([0,T]\times\mathbb{R}_{+})}\leq\left(\frac{\beta}{1-q}\right)\left(\frac{1-e^{-(\beta-\alpha)T}}{\beta-\alpha}\right).$
The next step consists in evaluating the norm of $\mathscr{T}^{n}$, for all $n\in\mathbb{N}^{*}$. By estimating the integral in a slightly different way, we have that
(10) $\begin{array}[]{c}\displaystyle\|\mathscr{T}^{n}g\|_{C_{b}([0,T]\times\mathbb{R}_{+})}\leq\left(\frac{\beta}{1-q}\right)^{n}\|g\|_{C_{b}([0,T]\times\mathbb{R}_{+})}\int_{0}^{t}\int_{0}^{t_{1}}\cdots\int_{0}^{t_{n-1}}\mathrm{d}t_{1}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}\leq\\[10.0pt] \displaystyle\left(\frac{\beta}{1-q}\right)^{n}\|g\|_{C_{b}([0,T]\times\mathbb{R}_{+})}\frac{T^{n}}{n!}.\end{array}$
We now claim that
$f(t,x)=\sum_{n=0}^{+\infty}\mathscr{T}^{n}F(f^{\text{in}})$
is a well defined function of $C_{b}([0,T]\times\mathbb{R}_{+})$ and that it solves the fixed-point problem (9). We indeed have that
$\begin{array}[]{c}\displaystyle\|f\|_{C_{b}([0,T]\times\mathbb{R}_{+})}\leq\displaystyle\sum_{n=0}^{+\infty}\|\mathscr{T}^{n}\|_{\mathcal{L}(C_{b}([0,T]\times\mathbb{R}_{+}))}\|F(f^{\text{in}})\|_{C_{b}([0,T]\times\mathbb{R}_{+})}\leq\\[10.0pt] \displaystyle\sum_{n=0}^{+\infty}\|f^{\text{in}}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}e^{|\alpha-\beta|T}\left(\frac{\beta T}{1-q}\right)^{n}\frac{1}{n!}=\|f^{\text{in}}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}e^{|\alpha-\beta|T}\exp\left(\frac{\beta}{1-q}T\right).\\ \end{array}$
Hence the series is normally convergent with respect to the natural norm of $C_{b}([0,T]\times\mathbb{R}_{+})$. Consequently, $f$ belongs to $C_{b}([0,T]\times\mathbb{R}_{+})$.
Moreover, $f$ solves (9) because
$f(t,x)=\sum_{n=0}^{+\infty}\mathscr{T}^{n}F(f^{\text{in}})=F(f^{\text{in}})+\sum_{n=1}^{+\infty}\mathscr{T}^{n}F(f^{\text{in}})=F(f^{\text{in}})+\mathscr{T}\sum_{n=0}^{+\infty}\mathscr{T}^{n}F(f^{\text{in}})=F(f^{\text{in}})+\mathscr{T}f.$
Uniqueness follows by contradiction. Suppose that there exist two distinct solutions $f_{1}$ and $f_{2}$ of (9).
Hence, their difference $\tilde{f}=f_{1}-f_{2}$ has finite and non-zero norm $\|\tilde{f}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}$. By linearity, $\tilde{f}$ solves
$\tilde{f}=F(f^{\text{in}})+\mathscr{T}\tilde{f},\qquad f^{\text{in}}=0.$
By iteration we have that
$\tilde{f}=\mathscr{T}\tilde{f}=\dots=\mathscr{T}^{n}\tilde{f}$
for all $n\in\mathbb{N}^{*}$ and, by passing to the norm, we have
$\|\tilde{f}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}=\|\mathscr{T}^{n}\tilde{f}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}$
for all $n\in\mathbb{N}^{*}$. Hence,
$\|\tilde{f}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}=0\text{ or }\|\mathscr{T}^{n}\tilde{f}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}$
The estimate given in (10) guarantees that $\|\mathscr{T^{n}}\tilde{f}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}\to 0$ when $n\to+\infty$, and hence this contradict the fact that $f_{1}$ and $f_{2}$ are distinct solutions of the fixed-point problem (9). ∎
5. Numerical tests
In this section we provide some numerical simulations which describe the behaviour of the model and compare it with the actual data.
5.1. Time evolution of the price sentiment
We study the problem
(11) $\left\{\begin{array}[]{l}\displaystyle\frac{\mathrm{d}}{{\rm d}t}s(t)=q\beta\big{[}X(t)(1+\Delta)-s(t)\big{]}\\[10.0pt] \displaystyle s(0)=s^{\text{in}}=\left(\int_{\mathbb{R}_{+}}xf^{\text{in}}(x)\mathrm{d}x\right)\\ \end{array}\right.$
with given $\Delta\in\mathbb{R}$, which has the explicit solution
$s(t)=\left(\int_{\mathbb{R}_{+}}xf^{\text{in}}(x)\mathrm{d}x\right)\left(\int_{\mathbb{R}_{+}}f^{\text{in}}(x)\mathrm{d}x\right)^{-1}\ e^{-q\beta t}+q\beta\int_{0}^{t}X(\theta)(1+\Delta)e^{q\beta(\theta-t)}\mathrm{d}\theta.$
We have developed a MATLAB script that implements a numerical solution for the differential equation (11) using a standard fourth-order Runge-Kutta method. We recall that the spot price $X(t)$ is derived from time-series data of the S&P 500 index, covering the period from May 26th, 2009, to November 20th, 2024.
To calibrate the model parameters and assess its accuracy, we divided the dataset into two subsets. The first subset, referred to as the training set, includes S&P 500 index data from May 26, 2009, to December 31, 2016. For the sentiment data $s_{\text{meas}}(t)$, we used the median of one-year analysts’ forecasts for the S&P 500 index over the same period. The training set consists of 1,987 data points.
These training data were used to optimize the model’s free parameters ($q$, $\beta$ and $\Delta$) using a combination of global and local optimization techniques. Specifically, a global optimizer was employed to explore the parameter space broadly, followed by a local optimizer for fine-tuning. The distance metric used in the optimization process was the standard Euclidean distance between the forecast vector $s(t)$ and the measured time series $s_{\text{meas}}(t)$. The optimized parameters were determined as $q=0.28$, $\beta=6.05$ and $\Delta=0.143$, which were subsequently used for forecasting.
It is important to note that, since 2009, the S&P 500 index has experienced significant growth interspersed with periods of volatility. Our dataset begins at the point when the index started recovering from the 2008 financial crisis. By 2013, the index had surpassed its pre-crisis levels and continued to grow steadily until 2019, despite challenges such as trade tensions and global economic uncertainties.
In 2020, the COVID-19 pandemic caused a sharp decline, with the index losing approximately 34% of its value by March. However, it recovered quickly due to government stimulus and strong performance in technology and healthcare sectors. In 2022, inflationary pressures and geopolitical tensions, including the war in Ukraine, triggered market corrections.
As a result, our training set is based on data from a period of market growth and does not include data from more volatile periods.
Using the optimized parameters, we simulated the period from January 3rd, 2017, to November 20th, 2024, and compared the results with the corresponding time series, consisting of 2,055 data points.
Figure 3 presents the results of our model compared to the median of one-year analysts’ forecasts for the forecasting set, which spans from January 3, 2017, to November 20, 2024, and to the cointegration model equilibrium which parameters are estimated over the same period as the kinetic model. For this simulation, the sentiment $s(0)$ was initialized using the value computed on January 3, 2017, $s(0)=2463.93$. It is clear that the kinetic model better captures the long-term dynamics compared to the cointegration model. Moreover, while the long-term equilibrium estimated by the cointegration model is much more volatile, closely mirroring the fluctuations of the S&\&&P 500 (due to $\alpha$ being near 1), the kinetic model does not suffer from the same volatility issues.
Figure 3. Forecasts $s_{\text{meas}}$ of the kinetic model (gray line), Forecasts $\exp{(\alpha\ln(X(t))+\beta)}$ of the cointegration model (dark blue line) and the time series of the median of one-year analysts’ forecasts for the S&P 500 index $s(t)$ (light blue line), from January 3rd, 2017, to November 20th, 2024.
Finally, in Figure 4, we plot the relative error between the forecasts generated by our model and the real data for the median of one-year analysts’ forecasts.
The statistical analysis of the relative errors reveals several insights into the robustness of the forecast. The mean relative error is equal to -0.00168, while the median is equal to -0.00217, indicating a slight systematic underestimation. This bias can be the consequence of the different market trends of the training set with respect to the forecasting set. The standard deviation of the errors, 0.00111, reflects a low level of dispersion, showing that the predictions are consistently close to the actual values.
The time series visualization of the relative errors shows how the errors fluctuate over time, with most values remaining close to zero. This consistency highlights the robustness of the forecast within a time horizon of approximately seven years and suggests that the model performs reliably across different periods.
One such event is the March 2020 market crash, triggered by the COVID-19 pandemic, which was followed by a rapid and sustained rally due to fiscal stimulus, monetary easing, and large-scale asset purchases by the Federal Reserve. Our model tended to overestimate forecasts before the crash and underestimate them afterward, reflecting a shift in market dynamics and a moderate increase in analysts’ overall optimism post-crash.
The model’s parameters were originally estimated using data up to 2016. To improve forecasting accuracy, a more dynamic calibration approach could be adopted, such as periodic updates of the parameters incorporating the entire historical dataset or a rolling window, for example of the past five years, to recalibrate at each time step.
Future research will further explore these methodologies, extending the analysis to other market indices to assess the generalisability of the findings.
Figure 4. Analysis of the relative error between the forecasts $s(t)$ of the model and the time series of the median of one-year analysts’ forecasts for the S&P 500 index $s_{\text{meas}}$, from January 3rd, 2017, to November 20th, 2024.
Acknowledgements. FS acknowledges the support of INdAM, GNFM group.
Conflicts of interest. The authors state that there is no conflict of interest.